OpenStudy (anonymous):

solve the following higher order differential equation? y^(6)-64y=0

OpenStudy (anonymous):

A fairly simple homogeneous equation. Solve the following characteristic equation: \[r^6-64=0\] Note that \(2^6=64\), so you have a difference of the form \(a^6-b^6\), which can be expressed as the difference of cubes \(\left(c^2\right)^3-\left(d^2\right)^3\). Basically, what I mean by all that, is that you can factor easily: \[r^6-64=0\\ \left(r^2\right)^3-\left(4\right)^3=0\\ \left(r^2-4\right)\left(\left(r^2\right)^2+4r^2+4^2\right)=0\\ (r+2)(r-2)\left(r^4+4r^2+16\right)=0\] From here, solve for \(r\).

OpenStudy (anonymous):

A minor detail I missed above: \[(r+2)(r-2)\left(r^4+4r^2+16\right)=0\] factors further to get \[(r+2)(r-2)(r^2+2r+4)(r^2-2r+4)=0\]

OpenStudy (anonymous):

Thank you so much,I'am check your solution it's right

OpenStudy (anonymous):

You're welcome!